# Majority voting is not good for heaven or hell, with mirrored performance

### v3

@@ Abstract
Within the ViSE (Voting in Stochastic Environment) model, we study the effectiveness of majority voting in various environments. By the pit of losses paradox, majority decisions in apparently hostile environments systematically reduce the capital of society. In such cases, the simple social decision rule of “rejecting all proposals without voting” outperforms majority. In this paper, we identify another pit of losses appearing in favorable environments. Here, the simple social decision rule of “accepting all proposals without voting” is superior to majority. We prove that under a version of simple majority called symmetrized majority and the antisymmetry of the voting body, the second pit of losses is a mirror image of the pit of losses in hostile environments and explain this phenomenon. Technically, we consider a voting society consisting of individualists whose strategy is supporting all proposals that increase their capital and a group (groups) whose members vote to increase the total group capital. According to the main result, the expected capital gain of each agent in the environment whose generator ${X}$ has mean $\mu>0$ exceeds by $\mu$ their expected capital gain under generator $-{X}$. This result extends to location families of generators with distributions symmetric about their mean. The mentioned result determines the symmetry of the difference between the expected capital gain under the symmetrized majority and that under the “basic” social decision rule that rejects/accepts all proposals in unfavorable/favorable environments.
@@
Voting paradoxes, Pit of losses, ViSE model, Majority voting, Symmetrized majority, Stochastic environment
@@ 1.2 The ViSE model
The main assumptions of the ViSE model (for more details, see e.g., [9]) are as follows. A society consists of $n$ agents (also called voters). Each agent is characterized by the current real-valued capital (debt, if it is negative), which can sometimes be interpreted as utility. In each step $t=1,\ldots,T,$ some proposal is put to the vote, and the agents vote, guided by their voting strategies, for or against it. In the framework of the ViSE model, a strategy is an algorithm following which an agent uses available information to decide whether to vote for or against the proposal under consideration. In general, such a strategy may involve stochastic elements. Each proposal is a vector of algebraic capital gains of all agents. Proposals approved through the established social decision rule are implemented: the agents’ capitals change according to the proposal. If the proposal is rejected, the capitals do not change.
Each proposal is generated stochastically: it is a realization ${\mathbf{{x}}}=({{x}_{1}},\ldots,{{x}_{n}})={{\mathbf{{X}}}(\omega)},\,\omega\in\Omega$ of a multivariate random variable ${\mathbf{{X}}}=({{X}_{1}},\ldots,{{X}_{n}})$ called a proposal generator. In this notation, $\omega$ is an elementary outcome, $\Omega$ is a sample space. In expressions like ${\mathbb{E}}({{\mathbf{{X}}}(\omega)}),$ where ${\mathbb{E}}$ is the mathematical expectation operator, $\omega$ is assumed to range over $\Omega,$ so that ${\mathbb{E}}({{\mathbf{{X}}}(\omega)})\equiv{\mathbb{E}}({\mathbf{{X}}}).$ We consider the case where the components of ${\mathbf{{X}}}$ are independent and identically distributed with a given mean $\mu$ and standard deviation $\sigma$. The scalar random variable ${X}$ corresponding to the components of ${\mathbf{{X}}}$ will be called the gain generator.
@@
The dynamics of agents’ capital in various environments can be analyzed to compare voting strategies and social decision rules in order to select the optimal ones in terms of maximizing appropriate criteria. Since in this version of the model the process is stationary, it can be described by one-step characteristics. An important one is the expected capital gain, ECG, of an agent with a certain strategy after implementing one collective decision of a given society in a given environment. The ECG (not to be confused with electrocardiogram) can be negative, zero, or positive.
@@ 1.3 Relation between the ViSE model and other models
or of some other agents with their own interests ([32, 34, 35, 39]).
@@
Proposals generated exogenously and having random effects were considered in ([38, 12, 13]) [^1] In [13], the importance of analyzing dynamic voting in a stochastic environment was emphasized and the scarcity of relevant work was noted. The authors mentioned publications since 2008; the ViSE model ([7], etc.) differs significantly from the Dziuda and Loeper model, but belongs to the same class., but the non-transferable nature of the individual utilities (i.e., measuring utilities in subjective scales rather than in a common “currency”) made it difficult to study cooperative and prosocial strategies within the corresponding models. A simulation study of majority voting by bots with a random agenda in a model with ideal points was conducted in [6].
@@
The studies of cooperation in dynamic voting in the game of pie sharing with transferable utilities started with [15]. An experimental study of cooperative solutions was performed in [14]. Interesting results on coalition dynamics and decision effectiveness in a model with discounting and bribes were obtained in [20]. The results of [9], where we looked for the optimal threshold of qualified majority, are comparable with the results of [24], where a stochastic model was studied for which, under certain conditions, simple majority was optimal. The problem of the optimal voting threshold was also considered in [10], where proposals were generated exogenously and the profitability of the first accepted proposal was evaluated; in that model, the unanimity rule would be optimal if there were no discounting.
@@
A significant difference between the ViSE model and legislative bargaining models (such as a number of the models mentioned above, see also [4, 42]) is that looking for equilibria is the main problem of bargaining models, while the structure of the proposal space in the ViSE model, with the individual capitals as the coordinates, does not allow any equilibria in standard problem settings. This is caused by the lack of Pareto optimal alternatives in the state space, as the basic ViSE model allows simultaneous benefit of all agents in any state. In such circumstances, the goal becomes the effectiveness of decisions made, i.e., utility maximization: for certain categories of agents or for the whole society. Thereby the ViSE model provides a means to study the effectiveness of agents’ selfishness, group cooperation, and various prosocial strategies, as well as the effectiveness of social decision rules.
@@
Thus, the ViSE model shifts the focus of research from finding equilibrium in bargaining to maximizing some aggregated utility, i.e., to the effectiveness under certain conditions.
@@
The analysis of effectiveness brings this approach closer to the one implemented in [22], where there is no dynamics, since a number of (exogenously generated) proposals are voted “in parallel”, but the acceptance of any proposal yields some change in individual utilities, and the characteristics of the resulting gain vector allow us to evaluate the effectiveness of the decision-making mechanism.
@@
It is worth noting that decision making within the ViSE model can be formalized as a general Markov decision process (see, e.g., [18, 41, 16]), where states are vectors of the current capital of the agents, and the possible social decision rules are as follows: vote by simple majority, vote using some other rules, accept the submitted proposal without voting, and reject the proposal without a vote. The last two rules can be treated as degenerate voting rules, namely, the threshold majorities with a negative threshold and with a threshold exceeding the number of voters, respectively. The probability (or density) of the transition between states $s$ and $s^{\prime}$ depends on the social decision rule and the probability (density) associated by the proposal generator with the proposal $s^{\prime}-s$. The society’s reward is the sum of the components of an accepted proposal, or zero if the proposal is rejected. This perspective is useful for interpreting research results within the ViSE model.
@@ 1.4 Pit of losses paradox
It is well known that agenda control allows, through a series of votes, to bring a voting society into a state worse than the initial state for all voters. Moreover, in most cases, the society can be brought to any state. This holds true for various models of sequential voting in a multidimensional policy space ([27, 28, 29, 32]) and is known as the so-called chaos theorems in social choice theory.
This is an important and paradoxical knowledge about social reality. It refutes the idea a la Adam Smith that a society composed of economic agents driven by self-interest will prosper through the invisible hand of democracy. No, through agenda control you can do anything you want to this society, even make it self-destruct. The majority’s gain may turn into the loss of all. There may be hope, however, that the issue is agenda manipulation, and by excluding it this effect can be eliminated.
The ViSE model leaves no room for a purposeful agenda: it is replaced by a favorable or unfavorable stochastic environment that generates proposals. However, a voting paradox of the above nature still arises. Namely, by [9, Proposition 1], if $\mu,$ the mean of a Gaussian gain generator ${X}$, is below a certain negative threshold (which is proportional to the standard deviation of ${X}$ and depends on the number of voters), then for a purely individualistic society and majority rule, the expected reward (one-step capital gain, ECG) of each agent is also negative. This is the pit of losses paradox illustrated in Fig. 1a. It means that sequential voting will bankrupt all participants in such a society.
@@
Figure 1: The expected one-step capital gain (ECG) of an agent in a voting society of 21 individualists: (a) in unfavorable environments $(\mu<0)$; (b) in some favorable environments $(\mu>0)$
This implies that in some unfavorable environments, majority rule is worse for each agent than rejecting all proposals without a vote, which saves capital. For gain generators distributed with longer tails than Gaussian generators, the pit of losses also exists, but it is shallower.
Thus, the essence of the paradox contained in the chaos theorems [^2] It should be noted, however, that sometimes decision making routines that encourage chaotic conflict are considered to be robust to bounded rationality and “promote effective search in an uncertain task environment” ([19]). Furthermore, chaos is not absolute as under a random agenda some areas are more attractive than others (see, e.g., [6]). is not agenda manipulation, but the combination of selfishness with majority rule. Methods to resolve the pit of losses paradox include cooperation ([8]), prosocial behavior ([43, 44]) that can be supported by income redistribution through taxes ([1]), and optimizing the voting threshold ([9, 26]; see Section 4 below).
@@ 1.5 Contributions of this study
The pit of losses paradox is characteristic of unfavorable environments. What happens in favorable environments? One might assume that the agent’s expected capital gain curve approaches the line $\mathrm{ECG}=\mu$ in the first quadrant from above as $\mu$ increases, i.e., that majority decisions provide some decreasing positive gain over what the environment offers.
However, somewhat surprisingly, this curve intersects the the line $\mathrm{ECG}=\mu$ (or $\mathrm{ECG}/\sigma=\mu/\sigma$, as shown in Fig. 1b). This means that under sufficiently favorable conditions, filtering proposals by majority vote is harmful. Furthermore, the abscissa of the intersection is equal in absolute value to the zero of the same function, marked with a black circle in Fig. 1a. In Sections 3 and 4 we study and explain this phenomenon in the more general case of a society consisting of individualists and a group (or groups).
Let us consider the following agent strategies. An individualist casts a vote (half a vote) for a proposal whenever this proposal increases (respectively, preserves) their own capital. The ‘half-vote’ outcome is provided to ensure the complementarity effect of opposite proposals in the case of discrete/discontinuous distributions. The voting strategy of a group member is to cast a vote (half a vote) for a proposal whenever this proposal increases (respectively, preserves) the total capital of the group.
@@
Lemma 1 below shows that under some conditions and any generator ${\mathbf{{X}}}$ with mean $\mu>0$, the expected capital gain of any agent exceeds that under generator $-{\mathbf{{X}}}$ by $\mu.$ In Theorem 3 and Remark 1, this result is applied to voting bodies consisting of individualists and group members. Theorem 4, an easy consequence of the previous one, states that the same difference of $\mu$ holds for generators ${\mathbf{{X}}}^{\mu}$ with mean $\mu$ and ${\mathbf{{X}}}^{-\mu}={\mathbf{{X}}}^{\mu}-2\mu,$ provided that ${\mathbf{{X}}}^{\mu}$ is symmetric about $\mu.$ In Section 4 we introduce the concept of expected relative capital gain (ERCG) and show that under the conditions of Theorem 4, the ERCG of any agent is an even function of $\mu,$ i.e., $\mbox{ERCG}(-\mu)=\mbox{ERCG}(\mu)$ for any $\mu$. This allows us to conclude that the pit of loses paradox extends to favorable environments, and the performance of the symmetric majority rule is symmetric about 00. This is why “majority voting is not good for heaven or hell, with mirrored performance” (see Section 4).
@@
2 Notation and problem formulation
We consider a society in which $\ell$ agents are individualists, while $g$ agents form a group, so that $g+\ell=n$ is the size of the society.
Let us say that any $n$ agents with their voting strategies and an adopted social decision rule form a voting body.
The social decision rule we consider is simple majority (a proposal is rejected/accepted if it receives no more/more than $n/2$ votes) modified by means of the random tie-breaking mechanism that accepts a proposal with probability $1/2$ when exactly $n/2$ votes are cast for it (cf., e.g., [17]). This rule will be called the symmetrized majority rule. When the gain generator is continuous and the number of agents is odd, this rule is equivalent to the simple majority rule, and we will use the latter name.
We say that a voting body is antisymmetric if for any proposal ${\mathbf{{x}}},$ the probabilities of ${\mathbf{{x}}}$ and $-{\mathbf{{x}}}$ being accepted by that voting body sum to $1$. An interpretation of this is that a voting body makes complementary stochastic decisions regarding any opposite proposals.
Given a voting body, let $n^{+}({\mathbf{{x}}})$ be the number of votes cast for proposal ${\mathbf{{x}}}$; let $I({\mathbf{{x}}})$ be the probability of accepting (and thus implementing) ${\mathbf{{x}}}$. If the social decision rule is not stochastic, then for any proposal ${\mathbf{{x}}},$ $I({\mathbf{{x}}})$ is either 00 or $1$
Let ${\mathbf{D}}_{\mathbf{{X}}}$ denote the random vector of capital gains that implements the collective decision on the proposal generated by ${\mathbf{{X}}}.$

### v4

@@ Abstract
Within the ViSE (Voting in Stochastic Environment) model, we study the effectiveness of majority voting in various environments. By the pit of losses paradox identified in previous work, majority decisions in apparently hostile environments tend to reduce the capital of society. In such cases, the simple social decision rule of “rejecting all proposals without voting” outperforms majority voting. In this paper, we identify another pit of losses appearing in favorable environments. Here, the simple social decision rule of “accepting all proposals without voting” is superior to majority voting. We prove that under a version of simple majority called symmetrized majority and the antisymmetry of the voting body, the second pit of losses is a mirror image of the pit of losses in hostile environments and explain this phenomenon. Technically, we consider a voting society consisting of individualists whose strategy is supporting all proposals that increase their capital and a group (groups) whose members vote to increase the wealth of their group. According to the main result, the expected capital gain of each agent in the environment whose generator ${X}$ has mean $\mu>0$ exceeds by $\mu$ their expected capital gain under generator $-{X}$. This result extends to location families of generators with distributions symmetric about their mean. The mentioned result determines the symmetry of the difference between the expected capital gain under the symmetrized majority and that under the “basic” social decision rule that rejects (resp. accepts) all proposals in unfavorable (resp. favorable) environments.
@@
Voting paradoxes, Pit of losses, ViSE model, Majority voting, Reversal symmetry, Stochastic environment, Random tie-breaking
@@ 1.2 The ViSE model
The main assumptions of the ViSE model (for more details, see e.g., [10]) are as follows. A society consists of $n$ agents (also called voters). Each agent is characterized by the current real-valued capital (debt, if it is negative), which can sometimes be interpreted as utility. In each step $t=1,\ldots,T,$ some proposal is put to the vote, and the agents vote, guided by their voting strategies. In the framework of the ViSE model, a strategy is an algorithm following which an agent uses available information to decide whether to vote for, against, or abstain from voting (a ‘neutral’ vote) on the current proposal. Such a strategy may involve stochastic elements.
A proposal is a vector of algebraic capital gains of all agents. It is generated stochastically as a realization ${\mathbf{{x}}}=({{x}_{1}},\ldots,{{x}_{n}})={{\mathbf{{X}}}(\omega)},\,\omega\in\Omega$ of a multivariate random variable ${\mathbf{{X}}}=({{X}_{1}},\ldots,{{X}_{n}})$ called a proposal generator. In this notation, $\omega$ is an elementary outcome, $\Omega$ is a sample space; ${\mathbf{{X}}}\!:\Omega\to{\mathbb{R}}^{n}$. We introduce $\Omega$ explicitly to distinguish it from another source of randomness arising from tie breaking.
We consider the case where the components of ${\mathbf{{X}}}$ are independent and identically distributed (i.i.d.) with a known finite mean $\mu.$ Let ${\mathcal{X}}^{\mu}$ be the set of such proposal generators. The scalar random variable ${X}{\color[rgb]{0,0,0}:\Omega\to{\mathbb{R}}}$ corresponding to a component of ${\mathbf{{X}}}$ will be called a gain generator.
@@
Proposals approved through the established social decision rule are implemented: the agents’ capitals change according to the proposal. If it is rejected, the capitals do not change. A social decision rule is an algorithm by which the decision is made to accept or reject a proposal ${\mathbf{{x}}}$ put to a vote. It may be stochastic and use votes cast by the agents, ${\mathbf{{x}}}$, and ${\mathbf{{X}}}$. A voting rule is a social decision rule that uses only agent votes.
The dynamics of agents’ capital in various environments can be analyzed to compare voting strategies and social decision rules in order to select the optimal ones in terms of maximizing appropriate criteria. Since in this version of the model the process is stationary, it can be described by one-step characteristics. An important one is the mathematical expectation of the capital gain (expected capital gain), ECG, of an agent with a certain strategy after implementing one social decision of a given society in a given environment. The ECG can be negative, zero, or positive.
@@ 1.3 Relation between the ViSE model and other models
or the prerogative of other agents with their own interests ([32, 34, 35, 39]).
@@
Proposals generated exogenously and having random effects were considered by [38] and [13, 14]111[14] emphasized the importance of analyzing dynamic voting in stochastic environments and noted the scarcity of relevant work. The ViSE model ([8], etc.) differs significantly from the Dziuda and Loeper model, but belongs to the same class., but the non-transferable nature of the individual utilities (i.e., measuring utilities in subjective scales rather than in a common ‘currency’) made it difficult to study cooperative and prosocial strategies within the models used. A simulation study of majority voting by bots with a random agenda in a model with ideal points was conducted by [6].
@@
Research on cooperation under dynamic voting in the pie-sharing game with transferable utilities was initiated by [16]. [15] performed an experimental study of cooperative solutions. [20] obtained interesting results on coalition dynamics and decision effectiveness in a model with discounting and bribes. The results of [10], where we looked for the optimal threshold of qualified majority, are comparable with the results of [24], where a stochastic model was studied for which, under certain conditions, simple majority was optimal. The problem of the optimal voting threshold was also considered by [11], where proposals were generated exogenously and the profitability of the first accepted proposal was evaluated; in that model, the unanimity rule would be optimal if there were no discounting.
@@
A significant difference between the ViSE model and legislative bargaining models (such as a number of the models mentioned above, see also [4, 42]) is that looking for equilibria is the main problem of bargaining models, while the structure of the proposal space in the ViSE model, with the individual capitals as the coordinates, does not allow any equilibria in standard problem settings. This is caused by the lack of Pareto optimal alternatives in the state space, as the basic ViSE model allows simultaneous benefit of all agents in any state. Therefore, the goal becomes the effectiveness of decisions made, i.e., utility maximization: for certain categories of agents or for the whole society. Thereby the ViSE model provides a means to study the effectiveness of agents’ selfishness, group cooperation, and various prosocial strategies, as well as the effectiveness of social decision rules.
@@
Thus, the ViSE model shifts the focus of research from finding equilibrium in bargaining to maximizing some aggregated utility, i.e., to the effectiveness under various conditions.
@@
The analysis of effectiveness brings this framework closer to the one studied by [22], where there is no dynamics, since a number of (exogenously generated) proposals are voted ‘in parallel’, but the acceptance of any proposal yields some change in individual utilities, and the characteristics of the resulting gain vector allow one to evaluate the effectiveness of the decision-making mechanism.
@@
It is worth noting that decision making within the ViSE model can be formalized as a general Markov decision process (see, e.g., [18, 41, 17]), where states are vectors of the current capital of the agents, and the possible social decision rules are as follows: vote by simple majority, vote using some other rules, accept the submitted proposal without voting, and reject the proposal without a vote. The last two rules can be treated as degenerate voting rules, namely, the threshold majorities with a negative threshold and with a threshold exceeding the number of voters, respectively. The probability (or density) of the transition between states $s$ and $s^{\prime}$ is determined by the social decision rule and the probability (density) associated by the proposal generator with the proposal $s^{\prime}-s$. The society’s reward is the sum of the components of an accepted proposal, or zero if the proposal is rejected. This perspective is useful for interpreting results obtained within the ViSE model (cf. Remark 1).
@@ 1.4 Pit of losses paradox
It is known since the 1970s that agenda control allows, through a series of votes conducted under simple majority rule, to bring a voting society into a state worse than the initial state for all voters. Moreover, in many cases, the society can be brought to any state. This is true for various models of sequential voting in a multidimensional policy space ([27, 28, 29, 32]) and is known as the so-called chaos theorems in social choice theory.
These theorems are an important and paradoxical knowledge about social reality. It refutes the idea à la Adam Smith that a society composed of economic agents driven by self-interest will prosper through the invisible hand of democracy. No, through agenda control you can do anything you want to this society, even make it self-destruct. The majority’s gain may turn into the loss of all. There may be hope, however, that the issue is agenda manipulation, and that by excluding it, the effect can be eliminated.
The ViSE model leaves no room for a purposeful agenda: it is replaced by a favorable or unfavorable stochastic environment that generates proposals. However, a voting paradox of the above nature still arises. Namely, by [10, Proposition 1], if $\mu,$ the mean of a Gaussian gain generator ${X}$, is below a certain negative threshold (which is proportional to the standard deviation of ${X}$ and depends on the number of voters), then for a purely individualistic society and majority rule, the expected reward (one-step capital gain, ECG) of each agent is also negative. This is the pit of losses paradox illustrated in Fig. 1a. It means that sequential voting will inevitably bankrupt all participants in such a society.
@@
Figure 1: The expected capital gain (ECG) of an agent in a voting society of 21 individualists: (a) in unfavorable environments $(\mu<0)$; (b) in some favorable environments $(\mu>0)$
This implies that in some unfavorable environments, majority rule is worse for each agent than rejecting all proposals without a vote, which saves capital. For some gain generators distributed with longer tails than Gaussian generators, the pit of losses also exists, but it is shallower.
Thus, the essence of the paradox contained in the chaos theorems [^2] It should be noted, however, that sometimes decision making routines that encourage chaotic conflict are considered to be robust to bounded rationality and “promote effective search in an uncertain task environment” ([19]). Furthermore, chaos is limited because with a random agenda, some areas are more attractive than others (see, e.g., [6]). is not agenda manipulation, but the combination of selfishness with majority rule. Methods to resolve the pit of losses paradox include cooperation ([9]), prosocial behavior ([43, 44]) that can be supported by income redistribution through taxes ([1]), and optimizing the voting threshold ([10, 26]; see Section 4 below).
@@ 1.5 Contributions of this study
The pit of losses paradox is characteristic of unfavorable environments (Fig. 1a). What happens in favorable environments? One might assume (cf. Fig. 1b) that the agent’s expected capital gain curve approaches the line $\mathrm{ECG}=\mu$ in the first quadrant from above as $\mu$ increases, i.e., that majority decisions provide some decreasing positive gain over what the environment offers.
However, somewhat surprisingly, this curve intersects the line $\mathrm{ECG}=\mu$ (or $\mathrm{ECG}/\sigma=\mu/\sigma$, as shown in Fig. 1b). This means that under sufficiently favorable conditions, filtering proposals by majority vote is harmful. Furthermore, the abscissa of the intersection is equal in absolute value to the zero of the same function, marked with a black circle in Fig. 1a. In Sections 3 and 4 we study and explain this phenomenon in the more general case of a society consisting of individualists and a group (or groups).
Let us consider the following agent strategies. An individualist casts a vote (half a vote) for a proposal whenever this proposal increases (respectively, preserves) their own capital. The ‘half a vote’ is a reasonable contribution in case of abstention. It ensures the complementarity effect of opposite proposals for discrete/discontinuous distributions in the results below. The voting strategy of a group member is to cast a vote (half a vote) for a proposal whenever this proposal increases (respectively, preserves) the total capital of the group this voter belongs to. Such a group can also be called a utilitarian group. Clearly, an individualist is a special case of a group.
@@
Lemma 1 below shows that under some conditions and any generator ${\mathbf{{X}}}$ with mean $\mu>0$, the expected capital gain of any agent exceeds that under generator $-{\mathbf{{X}}}$ by $\mu.$ In Corollary 6, this result is applied to voting bodies consisting of individualists and group members. Corollary 8, an easy consequence of the previous results, states that the same difference of $\mu$ holds for generators ${\mathbf{{X}}}^{\mu}$ with mean $\mu$ and ${\mathbf{{X}}}^{-\mu}={\mathbf{{X}}}^{\mu}-2\mu,$ provided that ${\mathbf{{X}}}^{\mu}$ is symmetric about $\mu.$ In Section 4 we introduce the concept of expected relative capital gain (ERCG). This value is the difference between the ECG under majority rule and that under the ‘basic’ social decision rule that rejects (resp. accepts) all proposals in unfavorable (resp. favorable) environments. We show that under the conditions of Corollary 8, the ERCG of any agent is an even function of $\mu,$ i.e., $\mbox{ERCG}(-\mu)=\mbox{ERCG}(\mu)$ for any $\mu$. This allows us to conclude that the pit of loses paradox extends to favorable environments, and the performance of the symmetrized majority rule is symmetric about 00. That is why “majority voting is not good for heaven or hell, with mirrored performance” (see Section 4).
In the above paragraph, we presented Corollaries 6 and 8, while Theorems 5 and 7 hold for the more general class of voting bodies formed by agents with ‘complementary’ strategies and a ‘complementary’ voting rule.
@@
2 Basic notation and problem formulation
We consider a society in which $\ell$ agents are individualists, while $g$ agents form a group, so that $g+\ell=n$ is the size of the society. Let us say that the set of agents with their voting strategies and an adopted social decision rule form a voting body.
The social decision rule we consider is the simple majority voting rule (a proposal is rejected/accepted if it receives no more/more than $n/2$ votes) modified by means of the random tie-breaking mechanism that accepts a proposal with probability $1/2$ when exactly $n/2$ votes are cast for it. This rule will be called the symmetrized majority rule. When the gain generator is continuous and the number of agents is odd, this rule is equivalent to the simple majority rule, and we will use the latter name.
Given a voting body, let $n^{+}({\mathbf{{x}}})$ be the number of votes cast for proposal ${\mathbf{{x}}}$; let $I({\mathbf{{x}}})$ be the probability of accepting (and thus implementing) ${\mathbf{{x}}}$. If the social decision rule and the agents’ strategies are not stochastic, then, for any ${\mathbf{{x}}},$ $I({\mathbf{{x}}})$ is either 00 or $1$
We say that a voting body is antisymmetric (or satisfies reversal symmetry; cf. [7]) if for any proposal ${\mathbf{{x}}},$ the probabilities of ${\mathbf{{x}}}$ and $-{\mathbf{{x}}}$ being accepted sum to $1$:
[…]
Such a voting body makes complementary stochastic decisions on opposite proposals.
Let ${\mathbf{D}}_{\mathbf{{X}}}$ denote the random vector of capital gains that implements the social decisions on the proposals generated by ${\mathbf{{X}}}.$ We note that ${\mathbf{D}}_{\mathbf{{X}}}$ is random due to its dependence on both ${\mathbf{{X}}}$ and the tie-breaking outcome. Therefore, the sample space of ${\mathbf{D}}_{\mathbf{{X}}}$ is the Cartesian product of $\Omega$ and the sample space of the tie-breaking rule.